Solve. (x + 3)2 = 6
Found 2 solutions by prabhjyot, Osran_Shri:
Answer by prabhjyot(165)
Answer by Osran_Shri(18)

(x + 3)2 = 6
if it is a simple equation then:
2x + 6 =6
2x = 12
x = 6
If it is a quadratic equation then, follw the solution as below:
(x+3)^2 = 6
x^2 + 3^2 + 2*x*3 = 6
x^2 + 6x = 6 - 6
x^2 + 6x = 0
This quadratic equation can be factored to solve for "x"
Use the principle of zero products,which says, if ab = 0, either a, b, or both must be equal to zero.
x ( x + 6 ) = 0
As per above principle,
x = 0 and x+6 = 0 ------> x = -6
Therefore x = 0, -6
B.Shridhar